# Transforming Code Repositories into MCP Services

### v2

@@ 3 Methodology: The Code2MCP Framework
12:   Review Agent: Analyze $\tau$ $\rightarrow$ generate correction plan $\delta$.
13:   Generation Agent: Re-synthesize MCP files using Code Report and correction plan $\delta$.
@@
17:Finalize Agent: Package service files, generate README, and create Pull Request.
@@
Initialization and Analysis. First, the Download Agent is responsible for cloning the specified repository, identified by its URL $u$, and creating an isolated local workspace. Subsequently, the Environment Agent begins its work by addressing environmental dependencies, one of the most common failure points in code conversion. It meticulously replicates the original code’s runtime environment by intelligently parsing dependency files or Dockerfiles. This step is the cornerstone for ensuring the reliability of all subsequent code generation and testing.
Once the environment is ready, the core challenge is identifying valuable, tool-worthy functionalities within a complex codebase. This is the task of the Analysis Agent. It moves beyond traditional static analysis by leveraging the DeepWiki tool for a deep semantic understanding of the code. By associating code entities with their intent from documentation and comments, the agent understands the code’s “purpose”, not just its “structure”. The final output is a detailed Code Report, which serves as the strategic blueprint for all subsequent stages.
Generation, Execution, and Self-Correction. After obtaining a clear conversion blueprint, the framework enters its core iterative loop, which is designed to tackle the challenge of translating abstract functionalities into concrete, executable, and correct MCP services.
The loop commences with the Generation Agent. It takes the Code Report from the Analysis Agent and uses the code generation capabilities of an LLM to intelligently abstract and encapsulate the identified core functionalities into MCP-compliant interfaces. This includes creating essential files like the tool interface definition and adapter file, which connect the original code to the MCP interface, along with a basic test suite.
Once the code is generated, the Run Agent immediately executes the test suite within the prepared environment. Its role is to verify the practical executability of the generated code. If the tests pass, the workflow exits the loop and proceeds to the finalization stage. However, if execution fails, the Run Agent captures a detailed error traceback, denoted as $\tau$, and passes it to the Review Agent, triggering the framework’s critical self-correction mechanism. The Review Agent acts as a “senior debugging engineer”, performing an in-depth error analysis. Its analysis goes beyond simply parsing the traceback $\tau$; it correlates the error message with the generated code, the original repository’s structure outlined in the Code Report, and the specific test case that failed. This contextual reasoning allows it to diagnose the root cause with high precision, identifying issues such as code logic error, a missing dependency, or an interface mismatch. After diagnosis, it formulates a correction plan, $\delta$. The resulting plan is a structured, actionable directive that specifies the exact file and code block to be modified and provides a clear description of the required change. This precise instruction is then fed back to the Generation Agent to re-synthesize the corrected code. This closed loop of “run-review-fix,” formed by the Run Agent and Review Agent, iterates until all issues are resolved and the generated code passes all tests. To prevent infinite loops, the process is bounded by a maximum of $B$ attempts. This robust self-correction mechanism is the key to Code2MCP’s end-to-end reliability.
Finalization and Delivery. Upon the successful completion of the core loop, the Finalize Agent handles the concluding tasks. It organizes and packages all the validated, functional MCP service files. To facilitate review and adoption by the original repository maintainers, it also automatically generates a clear README file explaining how to use the newly added MCP service. Finally, the Finalize Agent packages all generated artifacts into the clear directory structure shown in Figure 3. Operating as a fully automated contributor, it then uses the GitHub API to submit these additions as a new Pull Request to the original repository. This act completes the code-to-service transformation and builds a direct bridge connecting the open-source project to the growing MCP ecosystem.
4 Evaluation
4.1 Experimental Setup
Task. The evaluation task is to convert diverse GitHub repositories into validated MCP services, assessing both the efficiency in reducing conversion time and cost compared to manual effort, and the effectiveness demonstrated by the functional success of the generated tools.
Demonstrated Tools. As shown in Table 2, we demonstrate 6 representative open-source repositories to cover a spectrum of complexity and domains. The first group includes three widely-used utility libraries for simpler conversion tasks: TextBlob for natural language processing, dateutil for datetime parsing, and searchkick for intelligent search. These serve to demonstrate the significant speedup Code2MCP provides for standard conversion tasks. The second group includes three complex scientific libraries to test Code2MCP’s ability to handle domain-specific codebases: ESM for protein science, SymPy for symbolic mathematics, and Foam-Agent for computational fluid dynamics. Their primary purpose is to serve as the subjects of our in-depth case studies. These cases provide the qualitative evidence for its success on complex software. This capability is crucial for addressing the tool supply bottleneck.
Evaluation. We evaluate the effectiveness and generality of Code2MCP through a two-pronged approach: quantitative evaluation and in-depth case studies. The quantitative evaluation measures key performance metrics, such as conversion time and LLM token cost, against the estimated time for manual conversion to validate Code2MCP’s efficiency.
Table 2: Statistics on the time required for Code2MCP to convert code into MCP and the token consumption. The manual time is an average of estimates from three senior developers.
Repository Manual Time Code2MCP Speedup LLM Cost
(Est. hours) Time (min) Factor (Tokens)
sloria/TextBlob 1.0 – 1.5 12 $\sim$6x 24.7k
dateutil/dateutil 0.7 – 1.2 15 $\sim$4x 19.2k
ankane/searchkick 1.0 – 1.5 13 $\sim$6x 28.9k
facebookresearch/esm 4.0 – 5.0 50 $\sim$5x 58.5k
sympy/sympy 2.5 – 3.5 32 $\sim$6x 78.3k
csml-rpi/Foam-Agent 2.0 – 3.0 17 $\sim$8x 41.4k
4.2 Quantitative Evaluation
A core motivation of Code2MCP is to alleviate the significant manual effort of turning code repositories into MCP services. Therefore, we report the time and LLM tokens consumed by Code2MCP in Table 2 and compare this with the estimated manual effort for a human expert. The manual time is estimated by three senior developers familiar with the respective libraries and the MCP standard, and the average value is reported. The manual task includes code comprehension, code wrapping, protocol implementation, testing, and documentation. These tasks are tedious and time-consuming, posing a significant challenge even for experienced developers. From Table 2, we observe that the estimated manual time far exceeds that of Code2MCP. This disparity arises because the manual task is a complex endeavor, encompassing challenges like deep code comprehension and environment replication, which extend far beyond simple implementation and testing.
Having established the framework’s general efficiency, we now proceed to the case studies of the three scientific libraries. These studies on ESM, SymPy, and Foam-Agent will provide a qualitative analysis of the framework’s core capabilities.
4.3 Case Studies: Protein, Math, and Computational Fluid Dynamics
Bridging Protein Science with AI Agents. Advanced computational models, exemplified by AlphaFold (Jumper et al., 2021), have revolutionized structural biology with breakthroughs in protein structure prediction and sequence analysis. However, their practical application is hindered by steep learning curves and substantial engineering effort, creating a high barrier to entry for many scientists, especially those without deep computational backgrounds. For molecular biologists or drug discovery researchers focused on wet-lab experiments, writing Python scripts to analyze protein variants is often a significant distraction from their research. While AlphaFold is now present in the MCP ecosystem  [^3] https://github.com/augmented-nature/alphafold-mcp-server, ESM provides complementary capabilities such as zero-shot variant effect prediction to assess how mutations impact proteins, and it offers higher efficiency and accuracy. Thus, we select ESM as our conversion target. More broadly, scientific discovery depends on a diverse toolbox rather than a single universal solution, and the current scarcity of tool variety creates a critical “supply-side” bottleneck.
Historically, performing even foundational protein characterization with a library like ESM is a multi-step, technical endeavor. Users have to write custom scripts, navigate the library’s internal structure to call the correct functions, and manually parse the outputs. Visualizing the data often necessitates integrating separate third-party packages as well. This entire process is demanding of programming expertise, tedious, and error-prone, ultimately yielding key physicochemical properties.
In contrast, Code2MCP fully automates this entire process. The conversion itself is efficient, completed in about fifty minutes on a standard laptop. The resulting service hides the underlying technical complexity, making the tool accessible even to teams without programming expertise.
Code2MCP encapsulates the repository’s core capabilities into a powerful ESM MCP tool containing ten composable functions. These functions support a range of tasks, from foundational analysis to interpreting the structural effects of variants. As shown in Figure 4, when a user provides a protein sequence and a mutation site like “Q145G”, the AI agent invokes several functions from the tool. First, it calls the Analyze_sequence function to provide a baseline characterization, returning the protein’s key physicochemical properties such as molecular weight, amino acid composition, and hydrophobicity. Next, it calls the Predict_structure function to generate the protein’s 3D structure and uses a Predicted Aligned Error (PAE) plot to assess its reliability. Finally, the agent uses the Predict_variant_effect function. This function not only identifies the mutation’s location on the 3D structure and analyzes its potential impacts, such as disrupting local stability or altering ligand interactions, but it can also generate a 2D projection map to visually clarify the mutation’s spatial context and its proximity to key functional regions. By integrating the outputs from these functions, the agent provides a complete analysis, clearly outlining the entire process from sequence variation to structural change and, finally, to functional impact. This comprehensive analysis directly connects a protein variation to its potential structural and functional consequences, a critical step in fields like personalized medicine or protein engineering. Code2MCP enables an AI agent to manage this process end-to-end, delivering immediate and interpretable results without requiring code from the user. This automated conversion and simplified interaction reduces setup and runtime, allowing researchers to focus on scientific questions rather than code logistics.
Enhancing Mathematical Reliability in AI Agents. We next present a case study of Code2MCP in mathematics, where we create a comprehensive math MCP service by converting several core libraries, including SymPy. This case reflects a common challenge in science and engineering: LLMs are unreliable for precise computation. Their answers are difficult to verify, and minor errors can lead to incorrect conclusions. This forces agents to rely on their own numerical processing, making complex calculations a significant source of error and hallucination and undermining user trust in scientific and engineering applications. Meanwhile, powerful and validated codebases like SymPy are isolated from the modern AI ecosystem. Leveraging these libraries historically requires users to write custom Python scripts and manually translate mathematical problems using library’s specific functions and structures, a process that is both technically demanding and inefficient.
Using just the links to these math GitHub repositories, Code2MCP automatically generates a fully featured MCP service. This new service comprises 25 atomic functions covering a wide range of capabilities from symbolic algebra, calculus, and matrix operations to probability, statistics, and various mathematical transforms. Figure 5 shows a selection of these tools. For a specific problem, such as finding the indefinite integral of $y=\ln(x^{2}+1)$, an agent can invoke the newly added Integrate_expression function. Similarly, for a task like finding the volume of revolution for $y=\ln(x)\cdot\sin(x)$ from $x=1$ to $x=2\pi$, it uses the converted Calculate_volume function, producing an output that matches the result from an expert manually using the original library.
For more advanced research, an agent can leverage this suite of new functions for specialized, domain-specific calculations, extending its capabilities beyond simple arithmetic. This elevates the agent from a general-purpose assistant to a domain-aware collaborator capable of executing complex scientific workflows with verifiable precision. For example, a physicist can use the Laplace_transform function to analyze the stability of control systems or electrical circuits, while a digital signal processing expert can use the Z_transform function to analyze discrete-time signals for applications like designing audio filters. Furthermore, a financial analyst can employ the Fourier_transform to identify cyclical patterns in market data, thus unlocking new, specialized domains for AI agents by providing the validated functions they need.
Figure 5: The figure shows examples from the service’s mathematical toolkit, organized into three functional categories: “Value” for numerical calculus operations like calculating limits and areas; “Symbol” for symbolic algebra such as simplifying and integrating expressions; “Transform” for advanced operations like Fourier and Laplace transforms.
Case Study: Automating CFD Simulation for AI Agents. While the previous case studies focus on encapsulating individual functions, this section tackles the more profound challenge of automating multi-stage engineering workflows. Computational Fluid Dynamics (CFD) serves as a prime example of such a workflow. Traditionally, its interdependent steps require meticulous manual intervention. An engineer must write detailed scripts for tasks like geometry definition and manually edit complex text-based files to set solver parameters. This process is not only tedious and prone to syntax errors but also demands deep expertise in the repository’s internal structure. We selected Foam-Agent, a repository specifically for CFD simulations (Yue et al., 2025b). It is a framework designed to streamline operations within the OpenFOAM library (Jasak, 2009), capable of handling traditional tasks in CFD simulations such as mesh generation, solver configuration, and post-processing. However, its inherent execution logic follows a rigid, end-to-end sequential process, which lacks the flexibility required for dynamic adjustments or interrupt-resume scenarios in practical applications.
Code2MCP overcomes this rigidity by automatically decomposing Foam-Agent’s integrated simulation process into a set of independent atomic functions, such as Generate_mesh, Generate_simulate, and Visualize_velocity. This allows the agent to intelligently coordinate the entire complex simulation procedure through multi-turn dialogue with a user (Schick et al., 2024). This approach replaces tedious manual configuration file editing and transforms a command-line program into a flexible, interactive scientific partner. For example, the entire CFD simulation workflow, beginning with a general instruction such as “simulate flow over a cylinder” followed by the configuration of specific physical parameters, and concluding with the final request to “visualize the velocity field”, can be completed step-by-step through natural language dialogue between the user and the agent. As illustrated in Figure 6, this chained, interactive workflow:
  • Mesh Generation: The workflow commences with a user’s directive for a computational mesh, prompting the agent to invoke Generate_mesh to produce the geometric structure.
  • Simulation Execution: Following successful mesh generation, the user’s instruction defines the simulation’s physical parameters such as time steps and kinematic viscosity. The agent receives these parameters and uses the mesh as input for Generate_simulate automatically extracted by Code2MCP to configure and execute the simulation computation, abstracting away the complex syntax of solver control files, ensuring physical parameters are correctly formatted and validated. This function automates the creation of complex run scripts, such as Allrun or Slurm files, and manages the execution on either a local machine or a High-Performance Computing cluster, capturing log files for subsequent review.
  • Results Visualization: After the simulation converges, a user may request a visualization of a physical field. The agent then calls Visualize_velocity to render the simulation’s output. This function automatically processes the raw simulation data to generate insightful visual representations, such as 3D streamline plots. This process obviates the need for users to write complex post-processing scripts, enabling rapid analysis and interpretation of the simulation results.
Figure 6: AI agent orchestrates a CFD simulation pipeline by sequentially invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
5 Conclusion
Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework designed to automatically convert GitHub repositories into functional MCP services. The framework leverages a multi-agent workflow to handle core tasks of code comprehension, environment replication, and tool abstraction, while incorporating a “Run-Review-Fix” self-correction loop to enhance the reliability of the automation process. We demonstrate the framework’s feasibility and effectiveness by successfully converting several scientific computing libraries from fields including bioinformatics, symbolic mathematics, and computational fluid dynamics. This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, offering a promising solution to alleviate the MCP supply bottleneck and lower the barrier for its adoption.
Appendix A Appendix
A.1 Use of Large Language Models
A.2 Agent Roles and System Prompts
@@
Code Analysis System Prompt - Ingest repository signals (structure, import graph, README/docstrings, CLI entry points; DeepWiki if available) to identify stable public APIs suitable as MCP tools. - Prefer documented, side-effect-bounded surfaces; exclude tests, internals, notebooks, long demos unless clearly valuable and controllable. - Define crisp tool boundaries: explicit inputs/outputs, preconditions/postconditions, resource needs (CPU/GPU/memory/time), and I/O constraints. - Note minimal adapter needs (path normalization, dtype coercion, lazy imports) and hazards (network access, file mutation, global state). - Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds). - Also produce a case description: case name, case domain, case category, case solver. Ensure domain/category/solver belong to case_stats[’case_domain’/’case_category’/’case_solver’]. - Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions. Keep it actionable and minimal.
@@
Senior Software Engineer / Code Fixer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
@@
A.3 Example MCP Tool Implementations Generated by Code2MCP

### v3

@@ A.3 Detailed Conversion Pipeline and Algorithm
12:   Review Agent: Analyze traceback $\tau$ and generate correction plan $\delta$.
13:   Generation Agent: Re-synthesize MCP files using the Code Report and correction plan $\delta$.
@@
17:Finalize Agent: Package service files, generate README, and create a Pull Request.
@@
A.4 Agent Roles and System Prompts
@@
Code Analysis System Prompt - Ingest repository signals (structure, import graph, README/docstrings, CLI entry points; DeepWiki if available) to identify stable public APIs suitable as MCP tools. - Prefer documented, side-effect-bounded surfaces; exclude tests, internals, notebooks, long demos unless clearly valuable and controllable. - Define crisp tool boundaries: explicit inputs/outputs, preconditions/postconditions, resource needs (CPU/GPU/memory/time), and I/O constraints. - Note minimal adapter needs (path normalization, dtype coercion, lazy imports) and hazards (network access, file mutation, global state). - Summarize fragilities (optional deps, platform quirks) and propose guards (timeouts, argument validation, deterministic seeds). - Also produce a case description: case name, case domain, case category, and case solver, using a consistent taxonomy across repositories. - Output a compact plan for generation: candidate tools (name, brief description, inputs with types/defaults, outputs, idempotency, side effects) and environment assumptions. Keep it actionable and minimal.
@@
Senior Software Engineer This agent diagnoses failures and applies the smallest auditable change that restores correctness while preserving public contracts, deciding between direct fix and regeneration, using strict complete-file replacement, and avoiding multi-file edits or prose in outputs.
@@
A.5 Additional Protein Case Study: ESM
Appendix B Details of Evaluation
B.1 GPT-4 Template Baseline Prompt
In the GPT-4 template baseline, we use a single system prompt that instructs one GPT-4 model to generate a complete MCP service implementation for a given repository. The full prompt is given below.
GPT-4 Template Baseline System Prompt You are a single GPT-4 model acting as an autonomous MCP wrapper generator for a Python code repository. You receive: - A short task description. - The repository README (possibly truncated). - A list of key source files selected by static analysis, each with its file path and content. - A minimal description of the MCP protocol and the FastMCP library. Your goal is to produce a single, self-contained Python MCP server that exposes a small set of high-value tools from this repository. Requirements: 1. Design of tools - Identify 3–8 core, stable capabilities that are useful to expose as tools. - For each tool, choose a concise, descriptive name and a one-sentence natural-language description. - Specify explicit, typed parameters (no *args or **kwargs); include default values where reasonable. - Define clear, JSON-serializable return values. Avoid returning raw Python objects that cannot be serialized. 2. Implementation with FastMCP - Use the FastMCP library to implement the MCP server. - Implement a create_app() function that constructs and returns the MCP application. - Register each tool with FastMCP, binding it to the underlying repository functions or classes. - Prefer importing from public, documented APIs instead of private or test modules. - Use lazy imports for heavy optional dependencies when necessary, and return a helpful error message if a required package is missing. 3. Robustness and safety - Validate all inputs (types, value ranges, file existence) before calling library functions. - Handle exceptions inside each tool and return a standardized error object instead of raising uncaught exceptions. - The standardized response format for each tool is: { "success": bool, "result": any or null, "error": string or null }. - Avoid network access unless the repository explicitly requires it. - Restrict file I/O to the workspace; do not write to absolute paths outside the project. 4. Coding style - Write clean, idiomatic Python 3 code with clear function boundaries. - Add short docstrings for each tool describing its purpose, parameters, and return value. - Do not generate tests, examples, or documentation as part of the MCP server file. - Do not include Markdown fences such as ““‘; output only raw Python code. 5. Output format - Return a single Python file containing the complete MCP server implementation. - Do not include explanations, comments to the user, or multiple files; only the final server code that can be saved as e.g. mcp_service.py.
B.2 Evaluation Metrics and Bucketing Rules
This section defines the evaluation metrics and repository bucketing rules used in the main experiments.
Interaction steps. In the RepoMaster experiments, an interaction step is defined as a single high-level agent action that changes the external state of the system. Concretely, we count as one step any of the following: (1) invoking a tool (including Code2MCP-generated MCP tools and built-in RepoMaster tools), (2) editing source code files, or (3) executing the target program or tests. The total number of interaction steps for a task is the length of the resulting action sequence.
Task Set Description. In the integration experiments with RepoMaster and OpenAgents, we use a fixed set of code-understanding and tool-usage tasks that are primarily derived from public benchmarks and representative scenarios in prior work, with a small number of additional scientific-computing cases to ensure coverage of both general and scientific settings. At this stage, this task set remains a relatively small, research-oriented setup rather than a fully standardized large-scale benchmark, and constructing a more systematic MCP task benchmark is left as future work.
Task success rate and coverage. In the OpenAgents experiments, a task is considered successful if the agent produces a final answer that passes automatic correctness checks. Task success rate is the fraction of tasks in a group that are successful. Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct.
Repository buckets. To analyze the impact of dependency complexity and project style, we group repositories into four buckets based on static metadata:
  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt, environment.yml, or pyproject.toml, and primarily library-style public APIs.
  • Medium-dependency scientific libraries: between 11 and 30 Python dependencies with a library-style API and no substantial external system requirements.
  • High-dependency libraries: more than 30 Python dependencies or additional system-level requirements (e.g., MPI stacks, GPU-accelerated frameworks, or complex external services).
  • Script-heavy projects: projects whose primary functionality is exposed via CLI scripts, notebooks, or workflow-style pipelines rather than a stable library API; these are identified using repository structure, entry points, and README descriptions.
Appendix C Per-repository Results
Table 3: Repository identifiers, domains, and final outcomes.
ID Domain Repo Final_status Main_failure_type
B1 Biomedical facebookresearch/esm success
B2 Biomedical biocore/scikit-bio fail env_failure
B3 Biomedical Biopython/Biopython fail api_inference_error
B4 Biomedical pysam-developers/pysam success
B5 Biomedical deepchem/deepchem success
P1 Psychology psychopy/psychopy fail untoolable_repo
P2 Psychology pymc-devs/pymc success
P3 Psychology sahahn/BPt success
P4 Psychology mne-tools/mne-python fail api_inference_error
P5 Psychology neuropsychology/NeuroKit success
M1 Math sympy/sympy success
M2 Math scipy/scipy fail api_inference_error
M3 Math fredrik-johansson/mpmath success
M4 Math cvxpy/cvxpy success
M5 Math sagemath/sage success
Ea1 Earth Science obspy/obspy success
Ea2 Earth Science Unidata/MetPy fail mcp_spec_violation
Ea3 Earth Science pyproj4/pyproj fail repo_internal_bug
Ea4 Earth Science mapbox/rasterio success
Ea5 Earth Science geopandas/geopandas success
A1 Astronomy astropy/astropy success
A2 Astronomy sunpy/sunpy fail repo_internal_bug
A3 Astronomy lightkurve/lightkurve success
A4 Astronomy astroML/astroML success
A5 Astronomy astropy/astroquery success
C1 Chemistry rdkit/rdkit success
C2 Chemistry openbabel/openbabel fail env_failure
C3 Chemistry bjodah/chempy success
C4 Chemistry pyscf/pyscf fail api_inference_error
C5 Chemistry cclib/cclib success
Ph1 Physics csml-rpi/Foam-Agent success
Ph2 Physics PlasmaPy/PlasmaPy fail env_failure
Ph3 Physics pydy/pydy fail
Ph4 Physics PyAbel/PyAbel success
Ph5 Physics scikit-hep/scikit-hep fail untoolable_repo
S1 Social Science networkx/networkx success
S2 Social Science igraph/python-igraph success
S3 Social Science networkit/networkit fail env_failure
S4 Social Science snap-stanford/snap-python fail untoolable_repo
S5 Social Science datamade/dedupe success
L1 Linguistics nltk/nltk success
L2 Linguistics explosion/spaCy success
L3 Linguistics stanfordnlp/stanza success
L4 Linguistics RaRe-Technologies/gensim fail import_error
L5 Linguistics flairNLP/flair success
E1 Econometrics statsmodels/statsmodels success
E2 Econometrics alkaline-ml/pmdarima success
E3 Econometrics facebook/prophet fail env_failure
E4 Econometrics blue-yonder/tsfresh fail repo_internal_bug
E5 Econometrics pydata/xarray success
Table 4: Per-repository structural statistics (size, dependencies, and tests).
ID total_files total_size Size_bucket Dependency_complexity Has_tests Has_packaging
B1 471 33M Medium Medium Yes Yes
B2 874 9M Medium Medium Yes Yes
B3 1146 95M Medium High Yes Yes
B4 561 13M Medium Medium Yes Yes
B5 1411 136M Medium High Yes Yes
P1 3542 65M Large High Yes Yes
P2 371 17M Medium High Yes Yes
P3 2669 39M Large Medium Yes Yes
P4 1385 26M Medium High No No
P5 496 52M Medium Medium Yes Yes
M1 1968 29M Medium Low Yes Yes
M2 3027 81M Large High Yes Yes
M3 197 2M Small Low No Yes
M4 1079 38M Medium Medium Yes Yes
M5 5331 113M Large High Yes Yes
Ea1 2117 34M Large High Yes Yes
Ea2 595 99M Medium Medium Yes Yes
Ea3 144 1M Small Medium No Yes
Ea4 382 17M Medium Medium Yes Yes
Ea5 351 11M Medium Medium Yes Yes
A1 1879 24M Large High Yes Yes
A2 877 10M Medium Medium Yes Yes
A3 217 8M Small Medium No Yes
A4 192 1M Small Low Yes Yes
A5 960 23M Medium Medium Yes Yes
C1 5740 130M Large High Yes Yes
C2 10000 63M Large High Yes Yes
C3 213 1M Small Low No Yes
C4 2114 56M Medium High Yes Yes
C5 1508 80M Large Medium Yes Yes
Ph1 44 0.2M Small Medium Yes Yes
Ph2 498 16M Medium Medium Yes Yes
Ph3 263 8M Small Medium No Yes
Ph4 243 2M Small Low Yes Yes
Ph5 18 0.3M Small Low No Yes
S1 951 10M Medium Low Yes Yes
S2 248 3M Small Medium No Yes
S3 1168 22M Medium Medium Yes Yes
S4 678 14M Medium High Yes Yes
S5 97 1M Small Low No No
L1 496 8M Medium Low Yes Yes
L2 1683 12M Medium Medium Yes Yes
L3 574 5M Medium Medium Yes Yes
L4 638 55M Medium Medium Yes Yes
L5 383 5M Small Medium No Yes
E1 2091 39M Medium High Yes Yes
E2 225 1M Small Low No Yes
E3 282 14M Medium Medium Yes Yes
E4 135 10M Small Medium Yes Yes
E5 403 8M Medium Medium Yes Yes
Table 5: Per-repository environment and MCP conversion statistics.
ID Env_success MCP_success Num_tools Num_tools_passed First_run_status Num_fix_rounds
B1 Yes Yes 16 14 fail 1
B2 No No 0 0 fail 2
B3 Yes No 8 4 fail 2
B4 Yes Yes 10 9 success 0
B5 Yes Yes 14 13 fail 0
P1 No No 0 0 fail 1
P2 Yes Yes 9 8 success 0
P3 Yes Yes 11 10 success 0
P4 No No 13 12 fail 3
P5 Yes Yes 7 7 success 0
M1 Yes Yes 12 10 success 0
M2 Yes No 6 2 fail 2
M3 Yes Yes 8 7 success 0
M4 Yes Yes 10 9 success 0
M5 Yes Yes 15 13 fail 2
Ea1 Yes Yes 11 8 fail 2
Ea2 No No 5 1 fail 1
Ea3 No No 4 0 fail 2
Ea4 Yes Yes 9 8 success 0
Ea5 Yes Yes 7 7 success 1
A1 Yes Yes 8 5 fail 2
A2 No No 3 1 fail 1
A3 Yes Yes 10 9 success 0
A4 Yes Yes 6 6 success 1
A5 Yes Yes 8 8 success 0
C1 Yes Yes 18 16 fail 1
C2 No No 0 0 fail 2
C3 Yes Yes 7 7 success 0
C4 Yes No 12 10 fail 3
C5 Yes Yes 9 8 success 1
Ph1 Yes Yes 8 7 success 1
Ph2 No No 0 0 fail 2
Ph3 Yes No 6 6 fail 0
Ph4 Yes Yes 5 5 success 0
Ph5 No No 0 0 fail 1
S1 Yes Yes 10 9 success 0
S2 Yes Yes 8 7 fail 1
S3 No No 0 0 fail 2
S4 Yes No 0 0 fail 1
S5 Yes Yes 5 5 success 0
L1 Yes Yes 7 7 success 0
L2 Yes Yes 11 10 fail 2
L3 Yes Yes 9 8 success 1
L4 Yes No 0 0 fail 3
L5 Yes Yes 6 6 success 0
E1 Yes Yes 13 11 fail 1
E2 Yes Yes 7 7 success 0
E3 No No 0 0 fail 2
E4 No No 0 0 fail 1
E5 Yes Yes 8 7 success 1
Appendix D Example MCP Tool Implementations Generated by Code2MCP